### v1

# Squeezed Quasinormal Modes from Nonlinear Gravitational Effects

Abstract
We demonstrate that long-lived black hole quasi-normal modes become quantum mechanically squeezed due to nonlinear gravitational effects. We estimate the degree of squeezing for the fundamental mode of a Schwarzschild black hole quantitatively.
I Introduction
Consideration of the interaction of radiation with quantum detectors requires some assumption, implicit or explicit, about the quantum nature of the radiation. Classical treatment of the radiation, amounting to the replacement of its field strength by c-numbers, corresponds implicitly to the assumption that it is embedded in the quantum framework as a coherent state (coherent state hypothesis). Squeezing takes us outside the domain of coherent states, and brings in characteristically quantum-mechanical effects. Since nonlinearities at the source or in propagation generically induce squeezing, and general relativity is an inherently nonlinear theory, some amount of squeezing is likely to be present in all forms of gravitational radiation. It is a challenging task, however, to make quantitative estimates for realistic, complex sources. Here we address this challenge for the radiation emitted during the “ringdown” following black hole formation.
Black hole quasi-normal modes Regge and Wheeler (1957); Zerilli (1970); Vishveshwara (1970); Berti et al. (2009); Kokkotas and Schmidt (1999); Berti et al. (2025) describe this stage of radiation. Although their amplitude is relatively weak, they are (relatively) tractable mathematically. Indeed, earlier works have estimated the nonlinear-gravitational effects systematically in the observable ringdown signal Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Here we build upon these works to explore its implications for squeezing.
Expansion of Einstein’s field equations to the second order about a black hole spacetime,
$\tilde{g}_{\mu\nu}$ $\approx$ $g_{\mu\nu}+h_{\mu\nu}^{(1)}+h_{\mu\nu}^{(2)},$
[…]
suggests the following form for dynamical equations that describe the metric perturbations,
$G_{\mu\nu}^{(1)}[h^{(1)}]=0,~~\text{and}~~G_{\mu\nu}^{(1)}[h^{(2)}]=-G_{\mu\nu}^{(2)}[h^{(1)},h^{(1)}].$ (2)
Above $G_{\mu\nu}^{(i)},~i=1,2$ refers to the Einstein tensor $G_{\mu\nu}$ expanded to appropriate order Tomita (1974); Tomita and Tajima (1976); Cheung et al. (2023); Gleiser et al. (1996, 2000); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). In the Schwarzschild case, the first of the above equations leads to Regge-Wheeler Regge and Wheeler (1957), and Zerilli Zerilli (1970) equations for odd and even parity black hole quasi-normal modes, while the second equation leads to similar equations with a source term that is, up to derivatives, quadratic in the first order metric perturbations Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Recent works along this direction have noted that the coupled nature of wave equations at the second order imply higher harmonic generation of black hole quasinormal modes. There are good prospects to observe second harmonic generation sourced by lower order quasinormal modes, since the expected conversion ratio for lowest-order modes of non-rotating black holes is approximately, $\frac{|A_{4,4}(2\omega)|}{|A_{2,2}(\omega)^{2}|}\sim 0.15$ Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007). Perturbation theory for rotating (Kerr) spacetimes, notably developed by Teukolsky Teukolsky (1972) is more involved, but here too higher-harmonic quasinormal modes arise Ma and Yang (2024).
In what follows, we present a general framework inspired by these recent developments to describe potential probes for quantum effects in weakly damped (long-lived) quasinormal modes. The suggestion to use the nonlinearities in black hole ringdown as a probe for quantum gravity effects, based on the analogy to second harmonic generation in quantum optics, was also made in Ref. Guerreiro (2023). Here we provide a systematic approach. After deriving an effective interaction Hamiltonian, we estimate the degree of squeezing that it leads to quantitatively. In interesting limits we can relate our problem to a parametrically excited oscillator, a context in which squeezing has been quantified both theoretically and experimentally Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023). Our work complements recent discussions of how the coherent state hypothesis for gravitational radiation might be tested in practice with LIGO-type interferometers Parikh et al. (2020, 2021a, 2021b), and especially resonant mass detectors Tobar et al. (2024); Manikandan and Wilczek (2025a, ); Tobar et al. (2024); Manikandan and Wilczek (2025b); Shenderov et al. (2024).
Although we invoke most specifically black hole perturbation theory applied to the Schwarzschild background, the central ideas of our analysis are more widely applicable. Observed black hole events in LIGO correspond to the rotating (Kerr) black holes, where the lifetime of the fundamental quasinormal mode can be substantial, and in principle very large, becoming infinite in extremal limits Berti et al. (2009); Detweiler (1980); Yang et al. (2013a, b). While such modes are also difficult to excite, they provide limiting cases where our approach to estimating quantum effects, based on assuming long-lived modes, becomes exact. We defer a detailed analysis of such generalizations to future work.
II A simple theoretical model to describe quantum effects in quasinormal modes
Consider the following wave equation that naturally arise in the context of second-order quasinormal modes Cheung et al. (2023); Mitman et al. (2023); Nakano and Ioka (2007); Ioka and Nakano (2007):
$\left[\partial_{t}^{2}-\partial_{x}^{2}+\hat{V}\right]\psi=S.$ (3)
The homogeneous part of this equation corresponds to the homogeneous part of Eqns. (2), after separating out the angular variables using tensor harmonics. The potential $\hat{V}$ depends on whether we are considering scalar, electromagnetic (vector) or gravitational (tensor) perturbations. Our interest is in the gravitational, spin-two perturbations. In the non-rotating (Schwarzschild) context, the form of the equation or the potential can be broadly classified further in terms of the parity of the solutions. For the even parity (Zerilli Zerilli (1970)), solution change by $(-1)^{l}$ under the transformation $(\theta,\phi)\rightarrow(\theta-\pi,\phi+\pi)$, and for the odd parity perturbation (Regge-Wheeler Regge and Wheeler (1957)), the solution change by $(-1)^{l+1}$ under the transformation $(\theta,\phi)\rightarrow(\theta-\pi,\phi+\pi)$. Since the potential is also dependent on $l$, we have kept $\hat{V}$ as an operator such that $\left[\partial_{t}^{2}-\partial_{x}^{2}+\hat{V}\right]\psi_{lmn}=\left[\partial_{t}^{2}-\partial_{x}^{2}+V_{l}\right]\psi_{lmn}$ for stationary solutions $\psi_{lmn}$. After appropriate choice of coordinates it is possible to arrive at similar equations for rotating (Kerr) black holes, where however the potential $\hat{V}$ in the stationary, radial part of the wave equation also becomes frequency dependent Chandrasekhar and Detweiler (1997); Berti et al. (2025); Motohashi (2025).
The source term $S(x,t)$ relevant for second-order (quadratic) quasinormal modes typically includes, up to derivatives, a product of first-order quasinormal modes. The exact regularized form of the source term can be found elsewhere Ioka and Nakano (2007), which also respects the gauge invariant prescription from Garat and Price (2000). Also see Refs. Bucciotti et al. (2024); Redondo-Yuste et al. (2024). Given this, our first objective is to cast this general form of wave equation into a Hamiltonian formalism for the temporal modes, and subsequently estimate the degree of squeezing it generates on the fundamental mode. To estimate this, it is first instructive to simplify the wave equation to see how a quantum mechanical description can be applied to the early-time behavior of long-lived quasinormal modes.
We first consider the solution for the homogeneous part of the above equation using the method of separation of variables. The equation for the radial part looks like a time-independent Schrodinger equation,
$\psi(x,t)\rightarrow\tilde{\psi}(x,\omega):~~\left[-\partial_{x}^{2}+V(x,\omega)\right]\tilde{\psi}(x,\omega)=\omega^{2}\tilde{\psi}(x,\omega),$ (4)
where we have indicated that the potential $V$ could be frequency dependent, $V=V(x,\omega)$, to relate to the quasinormal modes of Kerr black holes Chandrasekhar and Detweiler (1997); Berti et al. (2025); Motohashi (2025). The solutions $\tilde{\psi}(x,\omega)$ typically satisfy the Siegert boundary conditions Siegert (1939): $\tilde{\psi}(x,\omega)\rightarrow e^{\pm i\omega x}$ as $x\rightarrow\pm\infty$ [in the choice of coordinates, $x\rightarrow$-$\infty$ corresponds to the event horizon]. As is well known, such boundary conditions for potentials relevant to the gravitational context imply that the mode frequencies $\omega=\{\omega_{n}\}$ [for given $l,m$, or generically $\omega=\{\omega_{lmn}\}$] are both complex and discrete, with negative imaginary parts. The corresponding solutions, or quasi-normal modes, are well characterized and understood Zerilli (1970); Regge and Wheeler (1957); Maggiore (2008); Berti et al. (2009); Vishveshwara (1970); Teukolsky (1972); Berti et al. (2025), using exact methods or with Wentzel-Kramers-Brillouin (WKB) type approximations Schutz and Will (1985). Once solutions to Eq. (3) are obtained, metric perturbations $h_{\mu\nu}^{(i)},~~i=1,2$ can be reconstructed by following a standardized prescription Berti et al. (2009); Motohashi (2025), while the temporal mode profiles are determined by the discrete quasinormal mode frequencies. In what follows, we suppress the $l,m$ indices often for simplicity, but unless specified otherwise, in the Schwarzschild context, they refer to the $l=4,m=4$ leading second-order quasinormal modes sourced by the $l=2,m=2$ modes.
For the homogeneous part, knowing the solutions of the form,
$h(t)\sim e^{-i\omega_{n}t},~~\text{where}~~\omega_{n}=\pm\omega_{R,n}-i\omega_{I,n},$ (5)
we can also recast the homogeneous part of equation (7) in the following form,
$\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})\left[\ddot{h}_{n}(t)+2\omega_{I,n}\dot{h}_{n}(t)+\Omega_{n}^{2}h_{n}(t)\right]=0.$ (6)
where $\Omega_{n}^{2}=\omega_{R,n}^{2}+\omega_{I,n}^{2}$ and $\omega_{I,n}\geq 0$ in our notation. Here it is evident that the temporal domain of the homogeneous part of the wave equation describes damped quantum harmonic oscillators, with damping coefficients $\gamma_{n}=2\omega_{n,I}$ and proper frequencies $\Omega_{n}=\sqrt{\omega_{R,n}^{2}+\omega_{I,n}^{2}}$. An insightful discussion of this correspondence can be found in Ref. Maggiore (2008).
Since it is the temporal modes that we quantize eventually, it is instructive to recast the original wave equation, Eq. (3), in its mode-expanded form where the spatial modes satisfy the quasinormal mode boundary conditions,
$\sum_{n}\tilde{\psi}_{n}(x,\omega_{n})\left[\ddot{h}_{n}^{(p)}(t)+\omega_{n}^{2}h_{n}^{(p)}(t)\right]=S(x,t),$ (7)
which is an abbreviation for,
$\sum_{n}\tilde{\psi}_{lmn}(x,\omega_{lmn})\left[\ddot{h}_{lmn}^{(p)}(t)+\omega_{lmn}^{2}h_{lmn}^{(p)}(t)\right]=S_{lm}(x,t),$ (8)
after silencing $l,m$ indices. The superscript $(p)$ indicates that we are interested in the particular solutions that account for the presence of a source term. In Eqns. (7)–(8), we have expanded the left-hand side in terms of the known spatial solutions of the homogeneous part that satisfy the quasinormal mode boundary conditions, in the spirit of treating the source term as effective force perturbations on the known homogeneous solutions. The temporal parts describe quantum harmonic oscillators, but with complex frequencies. We are interested in scenarios where the real part of the frequency dominates substantially, such that the quasinormal modes are long-lived. Practically, these are the modes that dominate the ringdown signal. The quantum behavior of these long-lived modes will then be effectively captured by the dynamics of their real frequency part.
II.1 Projection to quasinormal modes
To pick out the response of individual quasinormal modes, we need to consider some type of orthogonality relations for quasinormal modes. This is not straightforward, because the spatial modes of quasinormal modes are not square-integrable. Several alternatives that respect the “quasinormal” aspect of the modes have been considered in the literature Ching et al. (1995); Leung et al. (1997, 1998); Jaramillo et al. (2021); Green et al. (2023); London and Gurevich (2024); Motohashi (2025); Sauvan et al. (2022); Leaver (1986); Kristensen et al. (2020). They explore different approaches including include the use of Green’s functions Leaver (1986); Ching et al. (1995) and bi-orthogonal bases. Here one finds states $\{|\phi_{j}\rangle\}$ that are orthogonal to $\{|\tilde{\psi}_{n}\rangle\}$ such that the inner product $\langle\phi_{j}|\tilde{\psi}_{n}\rangle\propto\delta_{jn}$. Electromagnetic/optical analogies are discussed in Refs. Kristensen et al. (2020); Leung et al. (1994).
Our choice of inner product is motivated by treatment in Refs. Leung et al. (1994); Ching et al. (1995); Motohashi (2025); Leung et al. (1997). To define the inner product, we begin with the abbreviated notation $\psi_{j}=\tilde{\psi}(\omega_{j},x).$ It follows from Eq. (4) that,
$\psi_{j}\left[-\frac{\partial}{dx^{2}}+V(x,\omega_{n})\right]\psi_{n}=\omega_{n}^{2}\psi_{j}\psi_{n}~~\text{and}~~\psi_{n}\left[-\frac{\partial}{dx^{2}}+V(x,\omega_{j})\right]\psi_{j}=\omega_{j}^{2}\psi_{j}\psi_{n},$ (9)
where we have considered the generalized scenario where $V$ is $\omega$ dependent. Taking the difference of the two equations for $n\neq j$, we obtain,
$\psi_{j}\frac{\partial}{dx^{2}}\psi_{n}-\psi_{n}\frac{\partial}{dx^{2}}\psi_{j}=\left[U(x,\omega_{j})-U(x,\omega_{n})\right]\psi_{j}\psi_{n},~~U(x,\omega_{k})=\omega_{k}^{2}-V(x,\omega_{k}).$ (10)
Multiplying with $(\omega_{j}^{2}-\omega_{n}^{2})^{-1}$ on both sides and and integrating over a contour $\mathcal{C}$,
$\int_{\mathcal{C}}dx\frac{\left[U(x,\omega_{j})-U(x,\omega_{n})\right]}{(\omega_{j}^{2}-\omega_{n}^{2})}\psi_{j}\psi_{n}$ (11)
$=$ $\frac{1}{(\omega_{j}^{2}-\omega_{n}^{2})}\left\{\bigg{[}\psi_{j}\frac{\partial}{dx}\psi_{n}-\psi_{n}\frac{\partial}{dx}\psi_{j}\bigg{]}\bigg{|}_{\partial\mathcal{C}}-\int_{\mathcal{C}}dx\frac{\partial}{dx}\psi_{j}\frac{\partial}{dx}\psi_{n}+\int_{\mathcal{C}}dx\frac{\partial}{dx}\psi_{j}\frac{\partial}{dx}\psi_{n}\right\}.$
The last two terms cancel, and we have (for $n\neq j$),
$\int_{\mathcal{C}}dx~\frac{\left[U(x,\omega_{j})-U(x,\omega_{n})\right]}{(\omega_{j}^{2}-\omega_{n}^{2})}\psi_{k}\psi_{n}-\frac{1}{(\omega_{j}^{2}-\omega_{n}^{2})}\bigg{[}\psi_{j}\frac{\partial}{dx}\psi_{n}-\psi_{n}\frac{\partial}{dx}\psi_{j}\bigg{]}\bigg{|}_{\partial\mathcal{C}}=0.$ (12)
More formally, we can write as,
$\lim_{\omega_{k}\rightarrow\omega_{j}}\left\{\int_{\mathcal{C}}dx~\frac{\left[U(x,\omega_{k})-U(x,\omega_{n})\right]}{(\omega_{k}^{2}-\omega_{n}^{2})}\psi_{k}\psi_{n}-\frac{1}{(\omega_{k}^{2}-\omega_{n}^{2})}\bigg{[}\psi_{k}\frac{\partial}{dx}\psi_{n}-\psi_{n}\frac{\partial}{dx}\psi_{k}\bigg{]}\bigg{|}_{\partial\mathcal{C}}\right\}=0,$ (13)
when $n\neq j$. The limit $\omega_{k}\rightarrow\omega_{j}$ can be taken continuously along an appropriate contour in the complex plane of quasinormal mode frequencies, exploiting the circumstance that the frequencies $\{\omega_{n}\}$, although discrete for a given black hole, depend on continuously variable black hole parameters such as the black hole mass and spin. With the identification, $H=[-\partial_{x}^{2}+V(x,\omega)]$, we see from Eq. (4) that $H\psi_{j}=\omega_{j}^{2}\psi_{j}.$ This allows us to define the following operator (where the operations are performed in the order they appear),
$L(\psi_{j})[\cdot]$ $:=$ $\lim_{\omega_{k}\rightarrow\omega_{j}}\left\{\int_{\mathcal{C}}dx~\psi_{k}\left[U(x,\omega_{k})+\partial_{x}^{2}\right]\frac{1}{(\omega_{k}^{2}-H)}[\cdot]-\bigg{[}\psi_{k}\frac{\partial}{dx}-\bigg{(}\frac{\partial}{dx}\psi_{k}\bigg{)}\bigg{]}\frac{1}{(\omega_{k}^{2}-H)}[\cdot]\bigg{|}_{\partial\mathcal{C}}\right\},$
which obeys $L(\psi_{j})[\psi_{n}]=0$ when $j\neq n$ [follows from Eq. (13)]. Hence, the operator is very convenient to project out quasinormal modes other than $\psi_{j}$ from a linear expansion. Proceeding with the Schwarzschild scenario, we also note that
$L(\psi_{j})[\psi_{j}]$ $=$ $\lim_{\omega_{k}\rightarrow\omega_{j}}\int_{\mathcal{C}}dx~\frac{\left[U(x,\omega_{k})-U(x,\omega_{m})\right]}{(\omega_{k}^{2}-\omega_{j}^{2})}\psi_{k}\psi_{j}$ (15)
- $\lim_{\omega_{k}\rightarrow\omega_{j}}\bigg{[}\psi_{k}\frac{\partial}{dx}\psi_{j}-\bigg{(}\frac{\partial}{dx}\psi_{k}\bigg{)}\psi_{j}\bigg{]}\frac{1}{(\omega_{k}^{2}-\omega_{j}^{2})}\bigg{|}_{\partial\mathcal{C}}.$
The first term reduces to,
$\lim_{\omega_{k}\rightarrow\omega_{j}}\int_{\mathcal{C}}dx~\frac{\left[U(x,\omega_{k})-U(x,\omega_{j})\right]}{(\omega_{k}^{2}-\omega_{j}^{2})}\psi_{k}\psi_{j}=\int_{\mathcal{C}}dx~\frac{dU(x,\omega_{k})}{d\omega_{k}^{2}}\Bigg{|}_{\omega_{k}=\omega_{j}}\psi_{j}^{2}.$ (16)
To evaluate the second term, we may define [recalling the abbreviated notation $\psi_{j}=\tilde{\psi}(\omega_{j},x)$ for clarity],
$g(\omega_{k})=\tilde{\psi}(\omega_{k},x)\frac{\partial}{\partial x}\tilde{\psi}(\omega_{j},x)-\bigg{[}\frac{\partial}{\partial x}\tilde{\psi}(\omega_{k},x)\bigg{]}\tilde{\psi}(\omega_{j},x).$ (17)
Note that, $g(\omega_{j})=0$. This yields,
$\lim_{\omega_{k}\rightarrow\omega_{j}}\frac{g(\omega_{k})-g(\omega_{j})}{\omega_{k}^{2}-\omega_{j}^{2}}=\lim_{\omega_{k}\rightarrow\omega_{j}}\frac{g(\omega_{k})}{\omega_{k}^{2}-\omega_{j}^{2}}=\frac{\partial}{d\omega_{k}^{2}}g(\omega_{k})\bigg{|}_{\omega_{k}=\omega_{j}}.$ (18)
Hence we obtain,
$L(\psi_{j})[\psi_{j}]=\int_{\mathcal{C}}dx~\frac{dU(x,\omega_{k})}{d\omega_{k}^{2}}\Bigg{|}_{\omega_{k}=\omega_{j}}\psi_{j}^{2}-\frac{\partial}{d\omega_{k}^{2}}\bigg{[}\psi_{k}\frac{\partial}{dx}\psi_{j}-\bigg{(}\frac{\partial}{dx}\psi_{k}\bigg{)}\psi_{j}\bigg{]}_{\partial\mathcal{C}}\Bigg{|}_{\omega_{k}=\omega_{j}}:=\mathcal{N}_{j}^{2},$ (19)
where the norm $\mathcal{N}_{j}^{2}$ is complex in general Motohashi (2025); Leung et al. (1994); Ching et al. (1995). The action of $\mathcal{L}(\psi_{j})[\cdot]$ on the left-hand side of Eq. (7) only retains the coefficient of the $j^{\text{th}}$ mode in the expansion, with the normalization factor $\mathcal{N}_{j}^{2}$. The action of $\mathcal{L}(\psi_{j})[\cdot]$ on the right-hand side of Eq. (6) must therefore determine the driving term of the wave equation with appropriate pre-factors. Hence, we can write,
$\left[\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\right]=\mathcal{N}_{j}^{-2}\mathcal{L}(\psi_{j})[S(x,t)]=g_{j}(t).$ (20)
In the gravitational context, we know that the temporal part of the source contains the product of lower-order quasinormal modes, dominated by the fundamental ($\omega_{220}\equiv\omega_{0}$) mode Ioka and Nakano (2007). More generically, we can approximate, to the leading order, $g_{j}(t)\approx\sqrt{2}\Lambda_{j}[h_{j,0}(t)]^{2}$, where we have denoted $h_{j,0}$ as the fundamental mode having exactly half the frequency to that of the second harmonic mode $h_{j}^{(p)}(t)$. Recall that here the superscript $(p)$ denotes that $h_{j}^{(p)}(t)$ is the particular solution to Eq. (20). We have also absorbed any pre-factors to the effective parameter $\Lambda_{j}$, determined by Eq. (20). While these results apply to any of the modes, we restrict ourselves to the long-lived modes that have a substantial real frequency (propagating) part, which are also the ones likely to be observed. We thus obtain,
$\ddot{h}_{j}^{(p)}(t)+\omega_{j}^{2}h_{j}^{(p)}(t)\approx\sqrt{2}\Lambda_{j}[h_{j,0}(t)]^{2},$ (21)
where $h_{j,0}$ satisfies $\ddot{h}_{j,0}+\omega_{j,0}^{2}h_{j,0}=0$. The equation resembles the equation for a forced harmonic oscillator, except note that $\{\omega_{j}\}$ and potentially $\Lambda_{j}$, are complex. By considering long-lived modes, to the leading order, we may treat these parameters to be real and neglect the contributions from their imaginary parts. This only applies to lower-order quasinormal modes, as the higher-order ones are increasingly damped; however, the lower-order modes are also the ones of interest here.
Re-instating the notations in the Schwarzchild context as an example, to the leading order ($j=0$, $\Lambda_{0}\equiv\Lambda$), we have,
$\ddot{h}_{44}^{(p)}(t)+\omega_{440}^{2}h_{44}^{(p)}(t)\approx\sqrt{2}\Lambda[h_{220}(t)]^{2}~~\text{and}~~\ddot{h}_{kk0}(t)+\omega_{kk0}^{2}h_{kk0}(t)=0,~~k\in\{2,4\}.$ (22)
With $k=4$ in the second equation is of course the solution to the homogeneous part of the first equation. Also note that Eq. (22) is a slight generalization of the standard equation we often see in higher harmonic generation, since the source term has a frequency, $\omega_{220}\neq\omega_{440}$. Assuming a solution $h_{220}\approx a\cos{(\omega_{220}t)}$, for the fundamental mode, we obtain that the particular solution to the above equation has the following form,
$h_{44}^{(p)}(t)=\frac{a^{2}\Lambda}{\sqrt{2}\omega_{440}^{2}}+\frac{a^{2}\Lambda\cos(2\omega_{220}t)}{\sqrt{2}(\omega_{440}^{2}-4\omega_{220}^{2})}.$ (23)
Note that the temporal part of the particular solution oscillates at frequency $\omega_{44}^{(p)}=2\omega_{220}$, which is the second harmonic generated mode we are interested in. Anticipating quantization, we extract the Hamiltonian corresponding to the driving force,
$H_{I}\approx-\sqrt{2}\Lambda\left[\hat{h}_{220}(t)\right]^{2}\hat{h}_{44}^{(p)}(t).$ (24)
Returning to more general notation, let us consider the creation and annihilation operators for the field in the temporal domain $A_{\omega}$ for the fundamental mode and $A_{\omega}^{(p)}\equiv A_{\omega_{p}}\equiv A_{2\omega}$ for the second harmonic mode, we can rewrite the interaction Hamiltonian as
$H_{I}\approx-\frac{\hbar\lambda}{2}\left[(A_{\omega}^{\dagger})^{2}e^{-2i\omega_{0}t}+A_{\omega}^{2}e^{2i\omega_{0}t}+2A_{\omega}^{\dagger}A_{\omega}+1\right]\left(A_{\omega_{p}}e^{i\omega_{p}t}+A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right),$ (25)
where $\lambda\propto\Lambda\sqrt{\hbar}$ [see Appendix. A where $\lambda$ is estimated for gravitational radiation quantized within a cubical box of reduced volume $V=(c/\omega)^{3}$, where $\omega$ is the frequency of the radiation and $c$ is the speed of light.]. We can make a canonical transformation,
$\frac{1}{\sqrt{2}}\left(A_{\omega_{p}}e^{i\omega_{p}t}+A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right)\rightarrow\frac{i}{\sqrt{2}}\left(A_{\omega_{p}}e^{i\omega_{p}t}-A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right)$ (26)
and
$\frac{1}{\sqrt{2}i}\left(A_{\omega_{p}}e^{i\omega_{p}t}-A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right)\rightarrow\frac{1}{\sqrt{2}}\left(A_{\omega_{p}}e^{i\omega_{p}t}+A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right),$ (27)
so that the interaction Hamiltonian can be represented as,
$H_{I}\approx-i\frac{\hbar\lambda}{2}\left[(A_{\omega}^{\dagger})^{2}e^{-2i\omega_{0}t}+A_{\omega}^{2}e^{2i\omega_{0}t}+2A_{\omega}^{\dagger}A_{\omega}+1\right]\left(A_{\omega_{p}}e^{i\omega_{p}t}-A_{\omega_{p}}^{\dagger}e^{-i\omega_{p}t}\right).$ (28)
Noting that $\omega_{p}=2\omega_{0}$ (denoted as $2\omega$ in subscripts), we can make the rotating wave approximation to neglect the terms that oscillate at $\pm 2\omega_{0}$ and $\pm 4\omega_{0}$, and only keep terms that are energy conserving. Finally we obtain,
$H_{I}\approx-i\frac{\hbar\lambda}{2}\left[(A_{\omega}^{\dagger})^{2}A_{2\omega}-A_{\omega}^{2}A_{2\omega}^{\dagger}\right].$ (29)
Note that the fundamental quasinormal mode of the Schwarzschild spacetime is known to have a substantial (although not too large) quality factor $Q\approx 4.2$. While the quality factor remains fixed, the lifetime can increase with the mass of the black hole Berti et al. (2009). They could serve as sources of long-lived quasinormal modes, adequately described by the real-frequency (oscillatory) part over short time-scales relevant to resonant detection. The situation is especially favorable for rotating (Kerr) black holes, which are the ones also observed in LIGO. Here, the quality factors and lifetimes of the fundamental quasinormal modes can be substantially higher. For instance, the quality factors can increase with the black hole spin, approaching very large values near extremal conditions  Berti et al. (2009); Detweiler (1980); Hod (2008), even approaching the zero-damped limit Yang et al. (2013a, b). They correspond to extreme scenarios where our analyses become exact.
II.2 Approach to estimate $\lambda$ based on phenomenology
Above we have reduced the problem to a single effective parameter $\lambda$, potentially evaluated from the source term $S(x,t)$ using Eq. (20), or another appropriate choice of projection of the source term onto the quasinormal mode basis. It may be helpful to consdier a simpler phenomenological model that also allows for estimating the coupling $\lambda$ independently, and in principle, can be applied to extract the coupling from observed amplitudes in rather generic scenarios. To this end, we note that the amplitude $\frac{\lambda}{2}$ is determined by the fraction of energy dissipated via second harmonic generation, which numerical general relativity provides good estimates, that the amplitude of the second harmonic mode can be $\sim 10\%$ to that of the fundamental mode in some scenarios Cheung et al. (2023); Mitman et al. (2023). In fact, we can take the semi-classical limit to treat the fundamental mode in a coherent state $A_{\omega}|\alpha_{\omega}\rangle=\alpha_{\omega}|\alpha_{\omega}\rangle$, and approximate the interaction Hamiltonian as,
$H_{I}\approx-i\frac{\hbar\lambda}{2}\left[(\alpha_{\omega}^{*})^{2}A_{2\omega}-\alpha_{\omega}^{2}A_{2\omega}^{\dagger}\right].$ (30)
With these approximations, also neglecting the depletion of the fundamental mode, the effective dynamics reduces to a displacement operator on the second harmonic mode, which prepares a coherent state of the second harmonic mode with the amplitude,
$|\alpha_{2\omega}|\approx\left|\frac{\lambda}{2}\right||\alpha_{\omega}|^{2}t.$ (31)
It seems then appropriate that we could determine $\lambda$ at short times from simulations or actual data as,
$|\lambda|=\lim_{\Delta t\rightarrow 0}\frac{2}{\Delta t}\frac{|\alpha_{2\omega}(\Delta t)|-|\alpha_{2\omega}(0)|}{|\alpha_{\omega}|^{2}}.$ (32)
Numerical relativity can predict the classical gravitational wave amplitude ratios at early times, and assuming gravitational radiation quantized within a cubical box of reduced volume $V=(c/\omega)^{3}$, the classical gravitational wave amplitudes can be related to the coherent state amplitudes at early times as (see Appendix. A),
$\frac{h_{2\omega}}{h_{\omega}^{2}}\approx\sqrt{\frac{c^{5}}{8\pi G\hbar\omega^{2}}}\frac{|\alpha_{2\omega}|}{|\alpha_{\omega}|^{2}}\approx\sqrt{\frac{c^{5}}{8\pi G\hbar\omega^{2}}}\left|\frac{\lambda}{2}\right|t.$ (33)
We can use this amplitude ratio to estimate the degree of squeezing in the fundamental mode.
III Degree of Squeezing
The interaction Hamiltonian in Eq. (29) describes second harmonic generation, which is a well-known method in quantum optics to generate and probe sub-Poissonian states of the radiation field Mandel (1982). We can extend the analysis to the gravitational context as well, by considering quasinormal modes that are long-lived, and by restricting to timescales that are much smaller than the lifetime of the quasinormal modes. With this approximation, we can proceed to estimate the degree of squeezing generated in the ringdown phase. To this end, note that the Hamiltonian from Eq. (29) leads to the following set of equations that describe second harmonic generation,
$\frac{dA_{\omega}}{dt}=-\lambda A^{\dagger}_{\omega}A_{2\omega},~~\frac{dA_{2\omega}}{dt}=\frac{\lambda}{2}A_{\omega}^{2}.$ (34)
Now the analysis of squeezing can be carried out systematically, where we take $\lambda$ to be real for simplicity. With the fundamental mode initially in a coherent state $|\alpha_{\omega}\rangle$ and the second harmonic mode in its vacuum state, to leading order ($\lambda|\alpha_{\omega}|t\ll 1$), the covariance of the fundamental mode changes as,
$\langle(\Delta\hat{X}_{\omega})^{2}\rangle\approx\frac{1}{2}-\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4,~~\text{and}~~\langle(\Delta\hat{P}_{\omega})^{2}\rangle\approx\frac{1}{2}+\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4,$ (35)
that suggests that squeezing of the sub-Poissonian nature Mandel (1982). In Appendix B, we also map out a perturbatively exact analysis based on Refs . Ou (1994); Li and Kumar (1994); Suhara et al. (1996). For coherent state initial conditions for the fundamental mode and second harmonic mode in vacuum, where $\langle\hat{X}^{2}_{j\omega}(0)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle=\frac{1}{2}$, one obtains that,
$\langle\hat{X}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},~~j=1,2.$ (36)
Above, $\zeta=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t$ and,
$S_{\omega X}$ $=$ $(1-\zeta\tanh\zeta)^{2}\text{sech}^{2}\zeta+2\tanh^{2}\zeta\text{sech}^{2}\zeta,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)^{2}+\text{sech}^{4}\zeta,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\zeta+\zeta\text{sech}\zeta)^{2}+\text{sech}^{2}\zeta,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\zeta+(1-\zeta\tanh\zeta)^{2},$ (37)
which reduces to Eq. (35) at small $\zeta$. For $\zeta\gg 1$, one obtains that, $S_{\omega X}\rightarrow 4\zeta^{2}e^{-2\zeta}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\zeta}/8$, and $S_{2\omega P}\rightarrow\zeta^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996). This suggests substantial squeezing can in principle be generated through this mechanism, especially an exponential enhancement in the squeezing of the $\hat{P}$ quadrature of the fundamental mode.
To make some estimates relevant to our scenarios, note that we can approximate,
$|\zeta|=\frac{1}{\sqrt{2}}\lambda|\alpha_{\omega}(0)|t\approx\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right||\alpha_{\omega}(0)|,$ (38)
where we have used Eq. (33). For $h_{\omega}$ being the amplitude of the fundamental mode at the source, $|\alpha_{\omega}(0)|\approx h_{\omega}\sqrt{\frac{c^{5}}{32\pi G\hbar\omega^{2}}}$. Hence we obtain,
$|\zeta|=\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sqrt{\frac{c^{5}}{32\pi G\hbar\omega^{2}}}\approx\frac{1}{\sqrt{2}}\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|h_{\omega}\approx 0.1h_{\omega},$ (39)
since $\left|\frac{h_{2\omega}}{h_{\omega}^{2}}\right|\sim 0.15$. While the amplitude of the gravitational radiation is substantially small at our detectors, $h_{\omega}^{D}\approx 10^{-22}$, at the source it can be substantially high, and can be of the order of unity, suggesting that $|\zeta|\approx 0.1$. Assuming that the fundamental mode is in a coherent state, we are also in the regime of validity of Eq. (35),
$\langle(\Delta\hat{X}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}-\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1-|\zeta|^{2}\right)=0.495,~~\text{and}$
$\langle(\Delta\hat{P}_{\omega})^{2}\rangle$ $\approx$ $\frac{1}{2}+\lambda^{2}t^{2}|\alpha_{\omega}|^{2}/4=\frac{1}{2}\left(1+|\zeta|^{2}\right)=0.505.$ (40)
We thus estimate $S_{\omega X}\approx 0.99,~S_{\omega P}\approx 1.01$. This corresponds to a $1\%$ deviation in the noise of the fundamental mode due to second harmonic generation. While small, it is not minuscule. It serves as a sort of existence proof, showing the existence of a significant quantum effect in the tractable perturbative regime.
For reasons of concreteness and simplicity, we focused on a scenario where the fundamental mode begins in a coherent state. The merger dynamics may well prepare them in an acoherent quantum state, however. We defer a detailed analysis of this to future work. Also, we initialized the second harmonic mode in its vacuum state, but the merger dynamics could also excite the second harmonic mode. The resulting dynamics of the fundamental mode can then be understood in terms of a parametric oscillator, a natural scenario for describing quantum mechanical squeezing Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023). We discuss the mapping to a parametrically excited oscillator for the fundamental mode in an Appendix. C.
IV Discussion
We have estimated the degree of squeezing generated in long-lived quasinormal modes from gravitational self-interactions. The estimates are valid for the perturbative regime following the merger event. For this scenario, where only weak nonlinearities are involved, we find that higher harmonic generation of quasinormal modes leads to $\sim 1\%$ squeezing in the fundamental mode. As mentioned immediately above, our simplifying assumptions about initial conditions are not unlikely to be overly pessimistic. More generally, earlier stages of radiation close to merger, where nonlinearities are stronger, can be expected to yield larger effects. Recent work suggests possibilities for testing these ideas experimentally Manikandan and Wilczek (2025a, , b); Parikh et al. (2020, 2021b, 2021a); Tobar et al. (2024); Shenderov et al. (2024).
Several other astrophysical scenarios have also been suggested as potential candidates to generate quantum mechanical squeezed states of gravitational radiation. They include head-on collisions of black holes Lovas (2001) as well as the expansion of the universe, which could squeeze primordial gravitational waves Grishchuk and Sidorov (1989).
We conclude that observing characteristically quantum behavior in gravitational radiation is a challenging but potentially achievable goal.
V Acknowledgments
FW is supported by the U.S. Department of Energy under grant Contract Number DE-SC0012567 and by the Swedish Research Council under Contract No. 335-2014-7424. SKM was supported in part by the Swedish Research Council under Contract No. 335-2014-7424 and in part by the Wallenberg Initiative on Networks and Quantum Information (WINQ). We thank Maulik Parikh and Igor Pikovski for stimulating discussions around these subjects.
Appendix A Classical and quantum amplitudes of gravitational radiation
Quantized gravitational radiation within a cubical box of reduced volume $V=(c/\omega)^{3}$ in the linearized theory could be represented as (assuming a single polarization),
$h(t)$ $=$ $\frac{1}{\sqrt{2}}\sum_{\omega}\sqrt{\frac{16\pi G\hbar}{Vc^{2}\omega}}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t})=\frac{1}{\sqrt{2}}\sum_{\omega}\sqrt{\frac{16\pi G\hbar\omega^{2}}{c^{5}}}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t}).$ (41)
Assuming that the field is in a coherent state $|\alpha\rangle$, we can relate the number of gravitons to the intensity of the field (accounting for both polarizations),
$h_{0}^{2}\approx\frac{32\pi G\hbar\omega^{2}}{c^{5}}|\alpha_{\omega}|^{2}.$ (42)
Or alternatively, the number of gravitons,
$N\approx|\alpha|^{2}\approx\frac{h_{0}^{2}c^{5}}{32\pi G\hbar\omega^{2}}.$ (43)
Based on this, we can estimate the amplitude ratio,
$\frac{h_{2\omega}}{h_{\omega}^{2}}\approx\sqrt{\frac{c^{5}}{8\pi G\hbar\omega^{2}}}\frac{|\alpha_{2\omega}|}{|\alpha_{\omega}|^{2}}.$ (44)
The preceding calculations also allow us to estimate the effective coupling $\lambda$ as,
$H_{I}$ $\approx$ $-\sqrt{2}\Lambda\left[\hat{h}_{j,0}(t)\right]^{2}\hat{h}_{j}^{(p)}(t)$ (45)
$=$ $-\hbar\sqrt{\hbar}\Lambda\omega^{3}\left(\frac{16\pi G}{c^{5}}\right)^{3/2}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t})^{2}(A_{2\omega}e^{-i2\omega t}+A_{2\omega}^{\dagger}e^{i2\omega t})$
$=$ $-\frac{\hbar\lambda}{2}(A_{\omega}e^{-i\omega t}+A_{\omega}^{\dagger}e^{i\omega t})^{2}(A_{2\omega}e^{-i2\omega t}+A_{2\omega}^{\dagger}e^{i2\omega t}),$
where,
$\lambda=2\sqrt{\hbar}\Lambda\left(\frac{16\pi G\omega^{6}}{c^{5}}\right)^{3/2}.$ (46)
Appendix B Exact analysis of squeezing in second harmonic generation
We closely follow the analysis in Ref. Li and Kumar (1994); Ou (1994). Consider the dynamical equations that describe second harmonic generation,
$\frac{dA_{\omega}}{dt}=-\lambda A^{\dagger}_{\omega}A_{2\omega},~~\frac{dA_{2\omega}}{dt}=\frac{\lambda}{2}A_{\omega}^{2}.$ (47)
Define quantum operators about the classical mean-field solutions,
$a_{i\omega}=A_{i\omega}-\langle A_{i\omega}\rangle,~~i=1,2.$ (48)
They obey,
$\frac{d\langle A_{\omega}\rangle}{dt}=-\lambda\langle A_{\omega}\rangle^{*}\langle A_{2\omega}\rangle,~~\frac{d\langle A_{2\omega}\rangle}{dt}=\lambda\langle A_{\omega}\rangle^{2}.$ (49)
and
$\frac{da_{\omega}}{dt}=-\lambda(a^{\dagger}_{\omega}\langle A_{2\omega}\rangle+a_{2\omega}\langle A_{1\omega}\rangle^{*}),~~\frac{da_{2\omega}}{dt}=\lambda\langle A_{\omega}\rangle a_{\omega}.$ (50)
One can now define the following change of variables,
$u_{\omega}=\frac{|\langle A_{\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~u_{2\omega}=\frac{\sqrt{2}|\langle A_{2\omega}(t)\rangle|}{|\langle A_{\omega}(0)\rangle|},~~\text{and}~~\zeta=\frac{1}{\sqrt{2}}\lambda|\langle A_{\omega}(0)\rangle|t,$ (51)
in terms of which the mean-field equations are recast as,
$\frac{du_{\omega}}{d\zeta}=-u_{\omega}u_{2\omega},~~\frac{du_{2\omega}}{d\zeta}=u_{\omega}^{2},$ (52)
where perfect phase-matching is assumed, $\text{Arg}(\langle A_{2\omega}(t)\rangle)=\phi_{2\omega}(t)=2\text{Arg}(\langle A_{\omega}(t)\rangle)=2\phi_{\omega}(t).$ Hence, only the real amplitudes come into play. The system of equations admit solutions (assuming initial conditions $u_{\omega}(0)=1$, and $u_{2\omega}(0)=0$),
$u_{\omega}(\zeta)=\text{sech}\zeta,~~u_{2\omega}(\zeta)=\tanh\zeta.$ (53)
For the generalized position and momentum quadratures, $x_{j\omega}(\zeta)=\frac{1}{\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}+a^{\dagger}_{j\omega}e^{i\phi_{j\omega}})$ and $p_{j\omega}(\zeta)=\frac{1}{i\sqrt{2}}(a_{j\omega}e^{-i\phi_{j\omega}}-a^{\dagger}_{j\omega}e^{i\phi_{j\omega}}),$ one obtains equations,
$\frac{d\hat{X}_{\omega}}{d\zeta}$ $=$ $-\hat{X}_{\omega}\tanh\zeta-\sqrt{2}\hat{X}_{2\omega}\text{sech}\zeta,$
$\frac{d\hat{X}_{2\omega}}{d\zeta}$ $=$ $\sqrt{2}\hat{X}_{\omega}\text{sech}\zeta$
$\frac{d\hat{P}_{\omega}}{d\zeta}$ $=$ $\hat{P}_{\omega}\tanh\zeta-\sqrt{2}\hat{X}_{2\omega}\text{sech}\zeta,$
$\frac{d\hat{P}_{2\omega}}{d\zeta}$ $=$ $\sqrt{2}\hat{P}_{\omega}\text{sech}\zeta.$ (54)
Solutions are straightforward. One obtains,
$\hat{X}_{\omega}(\zeta)$ $=$ $(1-\zeta\tanh\zeta)\hat{X}_{\omega}(0)\text{sech}\zeta-\sqrt{2}\hat{X}_{2\omega}(0)\tanh\zeta\text{sech}\zeta,$
$\hat{X}_{2\omega}(\zeta)$ $=$ $\frac{1}{\sqrt{2}}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)\hat{X}_{\omega}(0)+\hat{X}_{2\omega}(0)\text{sech}^{2}\zeta,$
$\hat{P}_{\omega}(\zeta)$ $=$ $\hat{P}_{\omega}(0)\text{sech}\zeta-\frac{1}{\sqrt{2}}(\sinh\zeta+\zeta\text{sech}\zeta)\hat{P}_{2\omega}(0),$
$\hat{P}_{2\omega}(\zeta)$ $=$ $\sqrt{2}\hat{P}_{\omega}(0)\tanh\zeta+(1-\zeta\tanh\zeta)\hat{P}_{2\omega}(0).$ (55)
Based on this, one finds the covariance matrix elements (for initial conditions where $\langle\hat{X}^{2}_{j\omega}(0)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle=\frac{1}{2}$ and cross terms initially zero),
$\langle\hat{X}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{X}^{2}_{j\omega}(0)\rangle S_{j\omega X},~~\text{and}~~\langle\hat{P}^{2}_{j\omega}(\zeta)\rangle=\langle\hat{P}^{2}_{j\omega}(0)\rangle S_{j\omega P},$ (56)
where
$S_{\omega X}$ $=$ $(1-\zeta\tanh\zeta)^{2}\text{sech}^{2}\zeta+2\tanh^{2}\zeta\text{sech}^{2}\zeta,$
$S_{2\omega X}$ $=$ $\frac{1}{2}(\tanh\zeta+\zeta\text{sech}^{2}\zeta)^{2}+\text{sech}^{4}\zeta,$
$S_{\omega P}$ $=$ $\frac{1}{2}(\sinh\zeta+\zeta\text{sech}\zeta)^{2}+\text{sech}^{2}\zeta,$
$S_{2\omega P}$ $=$ $2\tanh^{2}\zeta+(1-\zeta\tanh\zeta)^{2}.$ (57)
For large $\zeta$, note that, $S_{\omega X}\rightarrow 4\zeta^{2}e^{-2\zeta}$, $S_{2\omega X}\rightarrow 1/2$, $S_{\omega P}\rightarrow e^{2\zeta}/8$, and $S_{2\omega P}\rightarrow\zeta^{2}$ Ou (1994); Li and Kumar (1994); Suhara et al. (1996), suggesting the exponential enhancement in the noise of the phase quadrature. While the predictions are limited by the validity of the linearized equations that, for instance, assume a highly intense coherent field on the fundamental mode to start with, we can take these results to indicate that substantial squeezing can, in principle, be generated via the second harmonic generation in the fundamental mode. However, for the scenarios we are interested in, we have also excluded the decay, or damping, of the quasinormal modes, so our predictions are relevant for weakly damped quasinormal modes that are long-lived, applicable at timescales much smaller than the lifetime of quasinormal modes. Hence, the relevant degree of squeezing can be estimated based on the limit of small $\zeta$, captured already by the perturbative expansion, Eq. (35) of the main text.
Appendix C Connections to a parametrically excited oscillator
The squeezing of the fundamental mode can also be understood in terms of a parametric oscillator if the second harmonic mode is populated to begin with. To see this, note that the following Lagrangian,
$L_{p}=\frac{1}{2}\left(\dot{h}_{44}^{(p)}\right)^{2}-\frac{\omega_{440}^{2}}{2}\left(h_{440}^{(p)}\right)^{2}+\sqrt{2}\Lambda h_{220}^{2}h_{440}^{(p)},$ (58)
yields our Eq. (22) for the second harmonic field,
$\ddot{h}_{44}^{(p)}+\omega_{440}^{2}h_{440}^{(p)}=\sqrt{2}\Lambda h_{220}^{2}.$ (59)
The Lagrangian for the free fundamental mode is given by,
$L_{0}=\frac{1}{2}\dot{h}_{220}^{2}-\frac{\omega_{220}^{2}}{2}h_{220}^{2},$ (60)
which in the limit $\Lambda=0$ generates the wave equation for the fundamental mode,
$\ddot{h}_{220}+\omega_{220}^{2}h_{220}=0.$ (61)
However, with nonzero $\Lambda$, variation of the total Lagrangian $L=L_{p}+L_{0}$ with respect to $\dot{h}_{220}$ yields,
$\ddot{h}_{220}+\left[\omega_{220}^{2}-2\sqrt{2}\Lambda h_{44}^{(p)}\right]h_{220}=0.$ (62)
We see that the mode $h_{440}^{(p)}$ modulates the frequency of the fundamental mode in a time-dependent manner. Since the temporal part of $h_{44}^{(p)}$ oscillates at frequency $\omega_{p}=2\omega_{220}$, we can approximate it as $h_{44}^{(p)}(t)\approx h_{0}^{(p)}\cos(2\omega_{220}t)$. With this, we note that the equation governing the fundamental mode is modified as,
$\ddot{h}_{220}+\left[\omega_{220}^{2}-2\epsilon\cos{(2\omega_{220}t)}\right]h_{220}=0,$ (63)
where $\epsilon=\sqrt{2}\Lambda h_{0}^{(p)}$. This is Mathieu’s equation in the standard form that describes a parametrically driven oscillator with $\omega_{p}=2\omega_{220}$. Such a drive naturally squeezes the fundamental mode as it can be understood as time-dependent modulations of the quantum harmonic potential, which has a squeezing effect Chaturvedi et al. (2002); Mallet et al. (2011); Murch et al. (2013); Wu et al. (1987); Adamyan et al. (2015); Manikandan and Qvarfort (2023).